OpenStudy (anonymous):

1 7x 13 - + - = - 2 10 20 i cant find x

OpenStudy (anonymous):

7x / 10 = (13/20) - (1/2) 7x / 10 = 3 / 20 14x / 20 = 3 / 20 So x = 3/14